b) Given the probability distribution below:
X 4 5 9 10
P(X) 1\2 1\6 1/6 1/6
Sketch the probability function and find the mean and standard deviation.
Answer

Answer by Boreal(15235)

Without the sketch
mean=4*0.5+(5+9+10)(1/6)
=2+4
=6
Variance=(4-6)^2*0.5+{(5-6)^2+(9-6)^2+(10-6)^2}*(1/6)
=2+4.33
sd=sqrt(6.33)=2.52